Let T:R2-R2 be multiplication by A. Determine whether T has an inverse; if so, find X2 5 81 -3 3 A- If inverse exists enter y1 and y2, otherwise enter NA for both. Click here to enter or edit your answer

show steps!
Answer #1

We know that {e1,e2} = {(1,0)T,(0,1)T} is the standard basis for R2. Further, T(e1) = A(1,0)T = (5,-3)T and T(e2) = A(0,1)T = (8,3)T. Hence, the standard matrix of T is M = [T(e1),T(e2)] =

5

8

-3

3

Further, since det(M)= 15+24 = 39 ? 0, hence M is invertible. Therefore, T-1 exists and T-1: R2? R2 is defined by T-1 (x1,x2)T = M-1 . (x1,x2)T.

Now, M-1 =

1/13

-8/39

1/13

5/39

so that T-1(x1,x2)T=M-1.(x1,x2)T=(x1/13-8x2/39, x1/13+5x2/39)T. Thus, y1= x1/13-8x2/39 and y2= x1/13+5x2/39.

The Chegg software is deleting all the mathematical symbols and replacing these by ?. Hence a screenshot is also attached.